the half life of Tl-201 is 73.1 hours. what's the decay constant?
Answer #1

Solution)

Given,

Half life, T=73.1 hrs=263160 s

Decay constant, lambda=?

We know,

T=0.693/lambda

So, lambda=0.693/263160 s= 2.63*10^-6 s^-1 (Ans)

==============